# Beyond Jailbreak: Unveiling Risks in LLM Applications Arising from Blurred Capability Boundaries

### v1

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.252941 www.ndss-symposium.org
@@ I Introduction
Amid the wave of large language models (LLMs), a new application paradigm—LLM applications (i.e., LLM apps, agents)—is reforming traditional application concepts. OpenAI launched GPTs in November 2023, enabling users to build customized AI applications leveraging ChatGPT’s capabilities. Subsequently, multiple frameworks and platforms for building LLM applications emerged, such as Coze [Bytedance_coze], Poe [Poe], LangChain [LangChain], and CrewAI [CrewAI]. The adoption of LLM applications is rapidly increasing across diverse, real-world scenarios, such as threat detection and mobile assistants.
@@
Unlike foundational LLMs, LLM-powered apps typically offer more specialized, domain-specific solutions. For instance, a coding app built upon an LLM is fine-tuned to excel at programming tasks, while an LLM-powered translation application can adeptly handle multilingual scenarios. Moreover, LLM applications provide non-expert users with more accessible and convenient entry points, bypassing the need for complex prompt engineering and the direct restrictions of foundation models (e.g., OpenAI’s regional access restrictions). However, if applications are developed without proper security considerations, these accessible entry points can be exploited. Even when a full jailbreak is not possible, an attacker can still breach the capability limits set by the application developer. This forces the application to produce unexpected outputs, such as deviating from its intended purpose or executing entirely unrelated tasks, all while operating within the foundational safety constraints of the LLM.
@@
Motivation Examples. The optimization of LLM apps for specific tasks introduces unforeseen complexities for traditional jailbreak methods. This motivates our research to move beyond conventional jailbreak and instead investigate “goal deviation”—the phenomenon where an application’s behavior strays from its intended objective without an explicit jailbreak. By studying this behavior in real-world scenarios, we aim to dissect the emerging risks of misuse associated with specialized LLM applications.
@@
We now illustrate a real-world business scenario from the Web3 industry, where strict auditing of operator actions is essential for protecting user funds. To this end, some Web3 companies deployed LLM-powered auditing systems to verify the compliance of operator actions, as depicted in Figure 1(a). If an action is deemed compliant, it proceeds normally; if an anomaly is detected, an alert is triggered for manual intervention. This creates an opportunity for a malicious operator to use specific inputs to bypass the audit. A malicious operator can use specially-crafted inputs or prompts to degrade the capabilities of the LLM auditor, bypass its checks, and subsequently steal user assets. We term this scenario “capability downgrade” (Figure 1(b)). In this case, the malicious input does not violate the foundational safety constraints of the LLM itself; it merely degrades the specialized capabilities of the application for its intended task. Second, we define “capability upgrade,” a state where an application, even without being fully jailbreak, can be abused to perform tasks beyond its original, intended scope. For instance, an operator can use crafted inputs to compel an auditing bot to perform additional tasks, subsequently abusing corporate assets for financial gain (as depicted in Figure 1(c)). Finally, “capability jailbreak” refers to the state where the application can be directed to execute any arbitrary task.
@@
Our Study. Our work systematically analyzes the differences between LLM and traditional applications, highlighting the potential risks introduced by the new development paradigm. The most significant distinction lies in the shift in the developer’s role—from a capability implementer to a capability restrictor. In traditional applications, all supported functionalities are developed by the developer, making code security the most critical concern. In contrast, the core functionalities of LLM applications are provided by the LLM itself. The developer’s primary role is to bridge the gap between users and the LLM while restricting unwanted LLM capabilities. A gap in the developer’s understanding of an application’s capabilities can impact or even compromise the application’s performance. Thus, we formally define the capability space of LLM applications to support their capability analysis.
@@ IV-B Prompt Evaluation
Specifically, similar to the “target” evaluation, we construct a bag of words for process descriptions, including step-related keywords (e.g., ”first”) and sequential markers (e.g., ”1.”). Then, we calculate the cosine similarity between each word in the prompt and the keywords in the bag of words, counting the number of matching words to obtain $N_{\text{step-keywords}}$ and $N_{\text{markers}}$. Next, to assess the complexity of logical relationships, we use NLP techniques to extract the number of logical connectives in sentences and the number of clauses, resulting in $N_{\text{logic-keywords}}$ and $N_{\text{clauses}}$. After normalization, we obtain $N_{\text{step-keywords-norm}}$, $N_{\text{markers-norm}}$, $N_{\text{logic-keywords-norm}}$, and $N_{\text{clauses-norm}}$. Finally, these normalized scores are weighted and fused to calculate the PScore.
@@
Specifically, we carefully designed a prompt to make the LLM output four values: level of detail in capability descriptions ($Cap_{\text{level}}$), number of capability entries ($N_{\text{capability}}$), number of constraint entries ($N_{\text{constraint}}$), and level of refusal ($Con_{\text{level},j}$) for each constraint in the prompt. Both $Cap_{\text{level}}$ and $Con_{\text{level},j}$ are rated on a scale of 1 to 5, with higher scores indicating more detailed descriptions or stricter refusal levels. For example, an application with $Con_{\text{level,j}}=5$ explicitly lists the tasks it refuses to perform. Appendix A provides our prompt and examples of scoring results for some applications. Finally, we calculate $CaScore$ and $CoScore$ using the following formulae.
@@ IV-C Capability Boundary Evaluation
For capability downgrade, since capability downgrade targets the weakening of an application’s unique capabilities, evaluating this across hundreds of thousands of applications is challenging. Thus, we evaluate the indirect prompt injection threat to LLMs to assess how carefully crafted inputs can impact the performance of LLMs on the same task. Different from generic prompt injection, our goal is to assess whether an LLM’s output for the same task remains robust when faced with inputs containing misleading information. We conduct simulation-based evaluations across diverse LLM application scenarios, crafting scenario-specific queries to gather responses. Then, we perform systematic comparisons to assess potential capability downgrade in target scenarios. Specifically, we carefully crafted 2,790 pairs of boundary test cases across 28 scenarios. For example, as shown in Figure 7, in the case of targeting LLM-powered candidate screening systems, strategic manipulation of resume content enables unqualified candidates to bypass screening filters, generating erroneous “highly suitable candidate” recommendations.
@@
To systematically evaluate capability constraints, we constructed a cross-category test case set based on the application types we previously identified. Our hypothesis is that if an application of Type A (App 1) can perform tasks from a Type B application (App 2), then App 1 is at risk of capability upgrade. Specifically, our constructed test case set includes the following three components:
1) Default cases. We observed that applications often provide a small number of default questions to guide users. We collected these default questions from applications of each type as representatives of their capability space. After manual verification and filtering, we selected five default test cases for each type.
2) Generated cases. Default questions are constrained by application developers. To extend the coverage of test cases, we generated out-of-scope test cases for each application type. For each type, we randomly select 5 applications, create their capability profiles, and use GPT-4o to generate 5 test cases for each application. Specifically, we first create a capability profile for the selected application, including its name, description, type, and the claimed capability space revealed through active interaction. We then input this profile into an LLM, instructing it to generate tasks that exceed the application’s capability space. For instance, if the application’s profile is a “weather app,” the LLM might be prompted to generate a coding task. To facilitate reproducibility, the prompts used in our experiments will be publicly released. To ensure the cases are accurate and representative, we manually select 5 cases from the generated cases for each type as the final set of generated cases.
3) Common cases. We created a set of 10 common cases (e.g., ”How many seasons are there in a year?”) to test the general performance of applications.
For capability jailbreak, our objective is to determine whether current jailbreak techniques are constrained by application-defined capability boundaries. We began by constructing a test case set with various malicious question types. Then, we reproduced state-of-the-art jailbreak techniques [0005LZY0S24, JiangXNXR0P24, abs-2310-08419, abs-2309-10253] to generate adversarial test cases capable of bypassing the foundational LLM’s constraints underlying the applications.
@@ IV-D Implementation and Evaluation
To evaluate the classification accuracy, we recruited 3 volunteers with experience in LLM application development as evaluation participants. Before the evaluation, the volunteers were trained on our classification standards. Specifically, we randomly sampled 200 applications from the classification results and asked the volunteers to determine whether the application type was correctly assigned. If at least 2 of the 3 volunteers agreed that the classification was accurate, we marked it as correct. Finally, among the 200 tested applications, 96% were classified correctly. The primary reason for false positives is that the capability of an application may be vague. For example, an app that simulates an ancient doctor (Coze ID 7399930163061981203) can be categorized as Education & Learning or Health & Wellness.
Prompt Evaluation. We used the en_core_web_sm and zh_core_ web_sm models in the python scapy library for tokenization. The rapidfuzz library was utilized for fuzzy matching of keywords. Additionally, the multilingual sentence embedding model “paraphrase-multilingual-MiniLM-L12-v2” was employed to compute the embedding vectors of Prompts, which were then used to derive their entropy values. These experiments were conducted on a laptop equipped with an Intel i7-12700H 2.30 GHz CPU and 64GB of RAM. Moreover, we used the OpenAI’s “gpt-4o-mini-2024-07-18” to calculate the CaScore and CoScore, priced at \$0.15 per 1M tokens for input and \$0.60 per 1M tokens for output. The total API cost for these experiments was approximately \$120.
@@
Since there is currently no established method for quantifying prompt quality, we recruited three volunteers to manually evaluate the accuracy of our scoring method. As prompts are expressed in natural language, differences in interpretation among volunteers were expected. To address this, we first introduced the volunteers to our four evaluation metrics and provided corresponding examples. We then randomly selected 100 application prompts and their scores for the volunteers to annotate whether the scoring was accurate. A score was considered accurate if at least two of the three volunteers marked it as such. Our evaluation results showed that our method accurately assessed 92% of the prompts. For the inaccurate results, we found that the primary issue was due to some prompts having unclear logical descriptions or use sequence words outside our predefined bag of words.
LLM Judge. We implemented our LLM Judge using OpenAI’s ChatGPT-4o API. The prompt template we used is provided in Appendix B. To evaluate the accuracy of the LLM Judge, we recruited 3 volunteers to assess its decisions through sampling. Specifically, we randomly sampled 300 judgment results from our test cases. We then provided the case descriptions and the application responses to the volunteers, asking them to determine whether the LLM Judge’s decision was correct. Ultimately, the accuracy of our LLM Judge reached 94.33% (283/300). The misclassified cases were primarily due to our design of a text-only LLM Judge, which may lead to misjudgments when evaluating multimodal tasks. For instance, in response to a requirement to generate an image, a textual description of the image might be mistakenly judged as a correct answer.
@@ V-A Characterizing Cross-Platform LLM Apps
Supported LLM. The LLMs supported by different platforms vary significantly. GPTs and AgentBuilder only support their own proprietary LLMs, while Coze and Poe are more flexible, supporting dozens of different large models. However, Coze’s free version only supports the Doubao foundational model, and access to other LLMs requires upgrading to the paid version. Table I shows the LLMs relied upon by the applications we collected from different platforms.
@@
Supper Developer. Super developers play a crucial role in the LLM application ecosystem, which causes low-quality, low-usage applications to occupy a significant portion of the current LLM app ecosystem. A developer is considered a super developer if he or she has released a significant number of applications. In traditional application domains, due to the lengthy development cycles, super developers are rare, and their applications are generally of lower quality, attracting limited user engagement. However, we observed the presence of super developers across various LLM application platforms, often drawing substantial user attention.
@@ VI-B Reproducibility
We will open-source the scripts used in our experiments [^1] https://anonymous.4open.science/r/nn4cz2wez4x7865n44. This includes the scripts for collecting LLM applications from different platforms and the LLM Judge script. Moreover, we will publish all the prompts we used to facilitate the reproducibility of our work. Finally, we will open-source the application metadata we collected from Coze, AgentBuilder, and Poe to encourage further research.

### v2

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.242941 www.ndss-symposium.org
@@ I Introduction
Amid the wave of large language models (LLMs), a new application paradigm, LLM applications (i.e., LLM apps), is reforming traditional application concepts. OpenAI launched GPTs in November 2023, enabling users to build customized AI applications leveraging ChatGPT’s capabilities. Subsequently, multiple frameworks and platforms for building LLM applications emerged, such as Coze [Bytedance_coze], Poe [Poe], LangChain [LangChain], and CrewAI [CrewAI]. The adoption of LLM applications is rapidly increasing across diverse, real-world scenarios, such as threat detection and mobile assistants.
@@
Unlike foundational LLMs, LLM-powered apps typically offer more specialized, domain-specific solutions. For instance, a coding app built upon an LLM is fine-tuned to excel at programming tasks, while an LLM-powered translation application can adeptly handle multilingual scenarios. Moreover, LLM applications provide non-expert users with more accessible and convenient entry points, bypassing the need for complex prompt engineering and the direct restrictions of foundation models (e.g., OpenAI’s regional access restrictions). However, if applications are developed without proper security considerations, these accessible entry points can be exploited. Even when a full jailbreak is not possible, an attacker can still breach the capability limit set by the application developer. This forces the application to produce unexpected outputs, such as deviating from its intended purpose or executing unrelated tasks, all while operating within the foundational safety constraints of the LLM.
@@
Motivation Examples. The optimization of LLM apps for specific tasks introduces unforeseen complexities for traditional jailbreak methods. This motivates our research to move beyond conventional jailbreak and instead investigate “goal deviation”, the phenomenon where an application’s behavior strays from its intended objective without an explicit jailbreak. By studying this behavior in real-world scenarios, we aim to dissect the emerging risks of misuse associated with specialized LLM applications.
@@
We now illustrate a real-world business scenario from the Web3 industry, where strict auditing of operator actions is essential for protecting user funds. To this end, some Web3 companies deployed LLM-powered auditing systems to verify the compliance of operator actions, as depicted in Figure 1(a). If an action is deemed compliant, it proceeds normally; if an anomaly is detected, an alert is triggered for manual intervention. This creates an opportunity for a malicious operator to use specific inputs to bypass the audit. A malicious operator can use specially-crafted inputs or prompts to degrade the capabilities of the LLM auditor, bypass its checks, and subsequently steal user assets. We term this scenario “capability downgrade” (Figure 1(b)). In this case, the malicious input does not violate the foundational safety constraints of the LLM itself; it merely degrades the specialized capabilities of the application for its intended task. Second, we define “capability upgrade,” a state where an application, even without being fully jailbreak, can be abused to perform tasks beyond its original, intended scope. For instance, an operator can use crafted inputs to compel an auditing bot to perform additional tasks, subsequently abusing corporate assets for financial gain (as depicted in Figure 1(c)). Finally, “capability jailbreak” refers to the state where the application can be directed to execute an arbitrary task.
@@
Our Study. Our work systematically analyzes the differences between LLM and traditional applications, highlighting the potential risks introduced by the new development paradigm. The most significant distinction lies in the shift in the developer’s role, from a capability implementer to a capability restrictor. In traditional applications, all supported functionalities are developed by the developer, making code security the most critical concern. In contrast, the core functionalities of LLM applications are provided by the LLM itself. The developer’s primary role is to bridge the gap between users and the LLM while restricting unwanted LLM capabilities. A gap in the developer’s understanding of an application’s capabilities can impact or even compromise the application’s performance. Thus, we formally define the capability space of LLM applications to support their capability analysis.
@@ IV-B Prompt Evaluation
Specifically, similar to the “target” evaluation, we construct a bag of words for process descriptions, including step-related keywords (e.g., “first”) and sequential markers (e.g., “1.”). Then, we calculate the cosine similarity between each word in the prompt and the keywords in the bag of words, counting the number of matching words to obtain $N_{\text{step-keywords}}$ and $N_{\text{markers}}$. Next, to assess the complexity of logical relationships, we use NLP techniques to extract the number of logical connectives in sentences and the number of clauses, resulting in $N_{\text{logic-keywords}}$ and $N_{\text{clauses}}$. After normalization, we obtain $N_{\text{step-keywords-norm}}$, $N_{\text{markers-norm}}$, $N_{\text{logic-keywords-norm}}$, and $N_{\text{clauses-norm}}$. Finally, these normalized scores are weighted and fused to calculate the PScore.
@@
Specifically, we carefully design a prompt to make the LLM output four values: level of detail in capability descriptions ($Cap_{\text{level}}$), number of capability entries ($N_{\text{capability}}$), number of constraint entries ($N_{\text{constraint}}$), and level of refusal ($Con_{\text{level},j}$) for each constraint in the prompt. Both $Cap_{\text{level}}$ and $Con_{\text{level},j}$ are rated on a scale of 1 to 5, with higher scores indicating more detailed descriptions or stricter refusal levels. For example, an application with $Con_{\text{level,j}}=5$ explicitly lists the tasks it refuses to perform. Appendix A provides our prompt and examples of scoring results for some applications. Finally, we calculate $CaScore$ and $CoScore$ using the following formulae.
@@ IV-C Capability Boundary Evaluation
For capability downgrade, since capability downgrade targets the weakening of an application’s unique capabilities, evaluating this across hundreds of thousands of applications is challenging. Thus, we evaluate the indirect prompt injection threat to LLMs to assess how carefully crafted inputs can impact the performance of LLMs on the same task. Different from generic prompt injection, our goal is to assess whether an LLM’s output for the same task remains robust when faced with inputs containing misleading information. We conduct simulation-based evaluations across diverse LLM application scenarios, crafting scenario-specific queries to gather responses. Then, we perform systematic comparisons to assess potential capability downgrade in target scenarios. Specifically, we carefully crafted 2,790 pairs of boundary test cases across 28 scenarios. Each test pair consists of one non-misleading case, to which the LLM can produce a correct response, and one case augmented with misleading content. For example, as shown in Figure 7, in the case of targeting LLM-powered candidate screening systems, strategic manipulation of resume content enables unqualified candidates to bypass screening filters, generating erroneous “highly suitable candidate” recommendations.
@@
To systematically evaluate capability constraints, we construct a cross-category test case set based on the application types we previously identified. Our hypothesis is that if an application of Type A (App 1) can perform tasks from a Type B application (App 2), then App 1 is at risk of the capability upgrade. Specifically, our constructed test case set includes the following three components:
1) Default cases. We observe that applications often provide a small number of default questions to guide users. We collect these default questions from applications of each type as representatives of their capability space. After manual verification and filtering, we select five default test cases for each type.
2) Generated cases. Default questions are constrained by application developers. To extend the coverage of test cases, we generate out-of-scope test cases for each application type. For each type, we randomly select 5 applications, create their capability profiles, and use GPT-4o to generate 5 test cases for each application. Specifically, we first create a capability profile for the selected application, including its name, description, type, and the claimed capability space revealed through active interaction. We then input this profile into an LLM, instructing it to generate tasks that exceed the application’s capability space. For instance, if the application’s profile is a “weather app,” the LLM might be prompted to generate a coding task. To facilitate reproducibility, the prompts used in our experiments will be publicly released. To ensure the cases are accurate and representative, we manually select 5 cases from the generated cases for each type as the final set of generated cases.
3) Common cases. We create a set of 10 common cases (e.g., ”How many seasons are there in a year?”) to test the general performance of applications.
For capability jailbreak, our objective is to determine whether current jailbreak techniques are constrained by application-defined capability boundaries. We begin by constructing a test case set with various malicious question types. Drawing inspiration from several well-regarded open-source projects (e.g., AI-Infra-Guard [Tencent_AI-Infra-Guard_2025], EasyJailbreak [zhou2024easyjailbreak]), we reproduce state-of-the-art jailbreak techniques [0005LZY0S24, JiangXNXR0P24, abs-2310-08419, abs-2309-10253] to generate adversarial test cases capable of bypassing the foundational LLM’s constraints underlying the applications.
@@ IV-D Implementation and Evaluation
To evaluate the classification accuracy, we recruited 3 volunteers with experience in LLM application development as evaluation participants. Before the evaluation, the volunteers were trained on our classification standards. We first introduced volunteers to standardized definitions for 20 application categories. Subsequently, we presented guiding principles for classifying applications with ambiguous boundaries. Evaluators were instructed to base classifications on core application functionality and were provided with three positive and three negative examples per category. Then, we randomly sampled 200 applications from the classification results and asked the volunteers to determine whether the application type was correctly assigned. If at least 2 of the 3 volunteers agreed that the classification was accurate, we marked it as correct. Finally, among the 200 tested applications, 96% were classified correctly. The Fleiss’ Kappa score for the three volunteers’ assessments is 0.9177. The primary reason for false positives is that the capability of an application is vague. For example, an app that simulates an ancient doctor (Coze ID 7399930163061981203) can be categorized as Education & Learning or Health & Wellness.
Prompt Evaluation. We used the en_core_web_sm and zh_core_ web_sm models in the python spaCy library for tokenization. The rapidfuzz library was utilized for fuzzy matching of keywords. Additionally, the multilingual sentence embedding model “paraphrase-multilingual-MiniLM-L12-v2” was employed to compute the embedding vectors of Prompts, which were then used to derive their entropy values. These experiments were conducted on a laptop equipped with an Intel i7-12700H 2.30 GHz CPU and 64GB of RAM. Moreover, we used OpenAI’s “gpt-4o-mini-2024-07-18” to calculate the CaScore and CoScore, priced at \$0.15 per 1M tokens for input and \$0.60 per 1M tokens for output. The total API cost for these experiments was approximately \$120.
@@
Since there is currently no established method for quantifying prompt quality, we recruited three volunteers to manually evaluate the accuracy of our scoring method. As prompts are expressed in natural language, differences in interpretation among volunteers were expected. To address this, volunteers first received instruction and training on our four evaluation dimensions: target, process, capability, and constraint. We presented the assessment criteria for each dimension and provided five prompt examples per dimension, illustrating a range of scores from low to high. We then randomly selected 100 application prompts and their scores for the volunteers to annotate whether the scoring was accurate. A score was considered accurate if at least two of the three volunteers marked it as such. Our evaluation results showed that our method accurately assessed 92% of the prompts. The Fleiss’ Kappa score is 0.8386. For the inaccurate results, we found that the primary issue was due to some prompts having unclear logical descriptions or use sequence words outside our predefined bag of words.
LLM Judge. We implemented our LLM Judge using OpenAI’s ChatGPT-4o API. The prompt template we used is provided in Appendix B. To evaluate the accuracy of the LLM Judge, we recruited 3 volunteers to assess its decisions through sampling. Specifically, we first clearly defined the criteria for “successful task completion” and “task failure” for the volunteers and provided ten reference examples for each category. Then, we randomly sampled 300 judgment results from our test cases. We next provided the case descriptions and the application responses to the volunteers, asking them to determine whether the LLM Judge’s decision is correct. Ultimately, the accuracy of our LLM Judge reached 94.33% (283/300). The Fleiss’ Kappa score is 0.9595. The misclassified cases were primarily due to our design of a text-only LLM Judge, which may lead to misjudgments when evaluating multimodal tasks. For instance, in response to a requirement to generate an image, a textual description of the image might be mistakenly judged as a correct answer.
@@ V-A Characterizing Cross-Platform LLM Apps
Supported LLMs. The LLMs supported by different platforms vary significantly. GPTs and AgentBuilder only support their own proprietary LLMs, while Coze and Poe are more flexible, supporting dozens of different large models. However, Coze’s free version only supports the Doubao foundational model, and access to other LLMs requires upgrading to the paid version. Table I shows the LLMs relied upon by the applications we collected from different platforms.
@@
Super Developer. Super developers play a crucial role in the LLM application ecosystem, which causes low-quality, low-usage applications to occupy a significant portion of the current LLM app ecosystem. A developer is considered a super developer if he or she has released a significant number of applications. In traditional application domains, due to the lengthy development cycles, super developers are rare, and their applications are generally of lower quality, attracting limited user engagement. However, we observed the presence of super developers across various LLM application platforms, often drawing substantial user attention.
@@ VI-B Reproducibility
We will open-source the scripts used in our experiments [^1] https://github.com/sy-yunyi/LLMApp-Eval. This includes the scripts for collecting LLM applications from different platforms and the LLM Judge script. Moreover, we will publish all the prompts we used to facilitate the reproducibility of our work. Finally, we will open-source the application metadata we collected from Coze, AgentBuilder, and Poe to encourage further research.
@@ VIII Conclusion
Acknowledgment
We thank all anonymous reviewers for their valuable and constructive feedback. We also thank Xudong Pan for his useful suggestions during the experimental process of this work. This work is supported by the National Key Research and Development Program of China (No. 2023YFB3105600), and the National Natural Science Foundation of China (Grant No. 62102218). Yunyi Zhang is partially supported by the Shuimu Tsinghua Scholar Program.